OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2. simplify (1/2A + 1/2B) and What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (kinggeorge):

Let's start with the first question. Simplify (1/2A + 1/2B). We know that A=1/2 and B=1/2. What is (1/2)*(1/2)?